# MLVU: A Comprehensive Benchmark for Multi-Task Long Video Understanding

### v2

@@ 5 Conclusion & Discussion
Limitations. While MLVU basically covers the major dimensions of long-video understanding, it can still be improved for even better comprehensiveness. For instance, there can be tasks related to high-resolution videos, or more specific tasks such as tracking and low-level processing. Our work will be a persistent effort, where new tasks and video types will be continually introduced.
@@ Appendix A Overview of Appendix
  • B: Collecting Details of our Universal Long Video Collection (ULVC).
  • C: Distribution of Video Durations for Each Task.
@@
  • E: Annotation Details of MLVU.
  • F: Details of Baselines and the Evaluation Process.
  • G: Explorations of Video Retrieval Augmented Generation.
  • H: More Experimental Results.
  • I: More Visualized Examples of MLVU.
  • J: Licensing, Hosting and Maintenance Plan.
  • K: Datasheet.
Appendix B Collecting Details of our Universal Long Video Collection (ULVC)
@@
Specifically, our ULVC consists of a diverse set of 757 long videos. This collection includes 168 movies sourced from the Movie101 dataset [52] and the MovieChat dataset [36], as well as 60 documentaries from the MovieChat dataset [36]. Additionally, it comprises 65 game videos from the MineDojo dataset [8], 200 surveillance videos from the UCF-Crime dataset [38], and 100 ego-centric videos from the Ego4D dataset [11]. Furthermore, our team independently collected 72 cartoons and 92 TV series to enrich the dataset.
It’s important to clarify that the quantity of videos in the ULVC does not directly correspond to the number of videos and questions in our MLVU benchmark, which are 1334 and 2593 respectively. For example, a two-hour movie from the ULVC might be utilized in its entirety for the Sub-Scene Captioning task, or it could be segmented into several approximately 10-minute clips for the Video Summary task, or even used as a background video for synthetic video generation. Moreover, a single video could be annotated with multiple questions simultaneously.
Appendix C Distribution of Video Durations for Each Task
Figure 4 illustrates the detailed video duration distribution for each task in our MLVU, featuring videos of varying lengths from three minutes to over 120 minutes.
Figure 4: Video duration distributions of different tasks in our MLVU.
@@ Appendix D Details of the MLVU Time-Ladder
Appendix E Annotation Details of MLVU
E.1 Topic Reasoning (TR).
@@ Appendix E Annotation Details of MLVU
E.4 Needle Question-Answering (NQA).
@@
E.6 Plot Question-Answering (PQA).
@@
E.7 Sub-Scene Captioning (SSC).
In the development process of the SSC task, we employed human annotators to generate both prompts and standard caption data. The specific guidelines provided to annotators are illustrated in Figure 10. Initially, the annotators identified a specific, easily referable sub-scene within a lengthy movie. Subsequently, they crafted a prompt replete with adequate clues to reference this scene, ensuring the uniqueness of these clues throughout the entire film. To prevent any leakage of information, the prompt was designed to exclude any character-specific names or objective hints, instead incorporating rich descriptive details to allude to the plot. Following this, the annotators produced a detailed caption for this sub-scene, and deconstructed the caption into multiple, non-redundant "scoring points" to facilitate quantitative assessment (the details of the evaluation metric can be found in Section F.3). More visualized examples of PQA task can be found in Figure 16.
@@
Appendix F Details of Baselines and the Evaluation Process
F.1 Baselines
In this section, we detail the primary baselines evaluated on our MLVU. For image-based MLLMs, most available models lack multi-image inference capabilities. Consequently, we select Otter-I, LLaVA-1.6, and InternVL, which offer official multi-image implementations. Additionally, we include three proprietary models—Claude-3-Opus, Qwen-VL-Max, and GPT-4 Turbo—that provide APIs for multi-image inference. For available models, we estimate the maximum input frames based on their maximum LLM context length. For Claude and Qwen, the supported maximum image numbers are approximately 20; therefore, we select 16 frames to ensure fair comparisons. Similarly, we choose 16 frames for GPT-4 Turbo to maintain consistency. In terms of video MLLMs, we adhere to the default settings for frame sampling strategies. For instance, VideoChat2 uniformly samples 16 frames, whereas LLaMA-Vid samples 1 frame per second. Specifically, the GPT-4o can support a maximum of approximately 500 images when each image’s resolution is set to 512$\times$512 pixels. Consequently, we opt for a sampling rate of 0.5 fps to accommodate the majority of our videos. Table 3 and Table 4 show the LLM-backbone as well as evaluated weight links used in each MLLM.
     Image MLLMs Short Video MLLMs Long Video MLLMs
  Model LLM Model LLM Model LLM
  Otter-I MPT-7B Otter-V LLaMA-7B Movie-LLM Vicuna-7B
LLaVA-1.6 Vicuna-7B mPLUG-Owl-V LLaMA-7B MovieChat Vicuna-7B
InternVL-1.5 Vicuna-13B VideoChat Vicuna-7B TimeChat LLaMA-2 7B
Claude-3-Opus Claude3 Video-ChatGPT Vicuna-7B LLaMA-VID Vicuna-7B
Qwen-VL-Max Qwen Video-LLaMA-2 Vicuna-13B MA-LMM Vicuna-7B
GPT-4 Turbo GPT-4 VideoChat2 Vicuna-7B MiniGPT4-Video Mistral-7B
Video-LLaVA Vicuna-7B GPT-4o GPT4
Table 3: LLM-Backbone in different MLLMs
     Image MLLMs Short Video MLLMs Long Video MLLMs
  Model Weight Model Weight Model Weight
  Otter-I Link Otter-V Link Movie-LLM Link
LLaVA-1.6 Link mPLUG-Owl-V Link MovieChat Link
InternVL-1.5 Link VideoChat Link TimeChat Link
Claude-3-Opus - Video-ChatGPT Link LLaMA-VID Link
Qwen-VL-Max - Video-LLaMA-2 Link MA-LMM Link
GPT-4 Turbo - VideoChat2 Link MiniGPT4-Video Link
Video-LLaVA Link GPT-4o -
Table 4: Link of evaluation weights. The specific version for “GPT-4 Turbo” and “GPT-4o” is “gpt-4-turbo-2024-04-09” and “gpt-4o-2024-05-13”.
F.2 Inference Detatils
@@ F.3 Evaluation Metrics
Appendix G Explorations of Video Retrieval Augmented Generation
@@
Rank Model Score
  Refer to caption GPT-4o 87.4
Refer to caption GPT-4 Turbo 79.5
Refer to caption InternVL-1.5 78.8
4 VideoChat2 74.6
5 Video-LLaVA 71.6
6 MiniGPT4-Video 70.9
7 Qwen-VL-Max 67.4
8 Claude-3-Opus 67.2
9 LLaVA 60.6
10 Video-LLaMA-2 54.5
(a) Anomaly Recognition
Rank Model Score
  Refer to caption GPT-4o 74.5
Refer to caption GPT-4 Turbo 68.0
Refer to caption InternVL-1.5 67.0
4 Qwen-VL-Max 63.5
5 Video-LLaVA 57.0
6 MiniGPT4-Video 52.5
7 VideoChat2 51.5
8 Claude-3-Opus 43.5
9 Video-LLaMA-2 41.5
10 LLaVA 41.0
(b) Topic Reasoning
Rank Model Score
  Refer to caption GPT-4o 4.90
Refer to caption GPT-4 Turbo 4.49
Refer to caption LLaMA-VID 3.22
4 InternVL-1.5 3.16
5 Claude-3-Opus 3.11
6 Movie-LLM 2.88
7 Qwen-VL-Max 2.71
8 MiniGPT4-Video 2.64
9 VideoChat2 2.57
10 TimeChat 2.54
(c) Video Summarization
Rank Model Score
  Refer to caption GPT-4o 64.8
Refer to caption Video-LLaVA 53.2
Refer to caption InternVL-1.5 52.7
4 MiniGPT4-Video 49.0
5 GPT-4 Turbo 45.9
6 LLaVA 43.1
7 MA-LMM 43.1
8 VideoChat2 42.0
9 Qwen-VL-Max 40.3
10 Video-ChatGPT 40.3
(d) Needle QA
Rank Model Score
  Refer to caption GPT-4o 57.1
Refer to caption MiniGPT4-Video 48.6
Refer to caption GPT-4 Turbo 47.4
4 VideoChat2 47.4
5 Video-LLaVA 45.2
6 InternVL-1.5 43.5
7 Video-ChatGPT 42.0
8 Qwen-VL-Max 40.9
9 Claude-3-Opus 40.2
10 MA-LMM 38.9
(e) Ego Reasoning
Rank Model Score
  Refer to caption GPT-4o 65.1
Refer to caption GPT-4 Turbo 60.6
Refer to caption InternVL-1.5 54.4
4 Video-LLaVA 48.4
5 Claude-3-Opus 47.8
6 MiniGPT4-Video 44.5
7 VideoChat2 43.8
8 Qwen-VL-Max 43.3
9 LLaVA 41.0
10 MA-LMM 35.8
(f) Plot QA
Rank Model Score
  Refer to caption GPT-4o 6.69
Refer to caption GPT-4 Turbo 6.21
Refer to caption Video-ChatGPT 5.48
4 mPLUG-Owl-V 5.31
5 Video-LLaVA 5.25
6 Video-LLaMA-2 5.22
7 LLaMA-VID 5.22
8 Qwen-VL-Max 5.21
9 VideoChat2 5.04
10 VideoChat 5.01
(g) Sub-Scene Caption.
Rank Model Score
  Refer to caption GPT-4o 56.7
Refer to caption InternVL-1.5 32.8
Refer to caption MovieChat 28.6
4 GPT-4V 26.5
5 LLaVA 25.5
6 Video-ChatGPT 25.1
7 MA-LMM 25.1
8 Qwen-VL-Max 25.0
9 TimeChat 24.7
10 VideoChat 24.3
(h) Action Order
Rank Model Score
  Refer to caption GPT-4o 46.3
Refer to caption Video-LLaVA 35.9
Refer to caption TimeChat 32.0
4 Video-ChatGPT 31.1
5 VideoChat2 29.6
6 VideoChat 28.6
7 LLaMA-VID 27.8
8 Otter-V 26.7
9 LLaVA 25.7
10 Video-LLaMA-2 25.7
(i) Action Count
Rank Model Score
  Refer to caption GPT-4o 64.6
Refer to caption InternVL-1.5 50.4
Refer to caption GPT-4 Turbo 49.2
4 Video-LLaVA 47.3
5 MiniGPT4-Video 44.5
6 VideoChat2 44.5
7 Qwen-VL-Max 42.2
8 LLaVA-1.6 39.3
9 Claude-3-Opus 36.5
10 MA-LMM 36.4
(j) M-Avg
Rank Model Score
  Refer to caption GPT-4o 5.80
Refer to caption GPT-4 Turbo 5.35
Refer to caption LLaMA-VID 4.22
4 InternVL-1.5 4.02
5 Qwen-VL-Max 3.96
6 Video-ChatGPT 3.90
7 mPLUG-Owl-V 3.84
8 Video-LLaVA 3.84
9 VideoChat2 3.81
10 Video-LLaMA-2 3.78
(k) G-Avg
Table 6: Leaderboard of different tasks in MLVU.
(until 2024/06/01)
Appendix H More Experimental Results
H.1 Leaderboard of each task in MLVU
Table 6 shows the leaderboard of each task in our MLVU.
H.2 Comprehensive evaluation results on MLVU Time-ladder
Table 7, Table 8 and Table 9 show the experimental results on MLVU Time-ladder within 3 minutes, 6 minutes and 10 minutes. As anomaly recognition and topic reasoning are global tasks, thus, we can not annotate them in MLVU Time-ladder as the way provided in Appendix D.
  Model Needle QA Ego Rea. Plot QA Action Or. Action Co. Sub-Scene Cap. Video Sum. M-Avg
  Otter-I 25.0 25.0 25.4 25.0 25.0 4.04 2.59 25.1
LLaVA-1.6 48.0 43.0 45.2 29.5 28.5 4.48 2.40 38.8
InternVL-1.5 70.5 47.5 60.9 19.5 28.5 5.42 3.30 45.3
Claude-3-Opus 34.5 42.5 38.1 15.5 22.0 4.12 3.35 30.5
Qwen-VL-Max 49.5 42.2 50.3 27.7 16.3 5.58 2.95 37.2
Otter-V 30.5 32.0 18.3 21.0 29.5 4.32 2.63 26.3
mPLUG-Owl-V 31.0 29.5 22.3 29.0 24.0 5.35 2.68 27.2
Video-LLaMA-2 55.0 35.0 36.5 22.5 22.0 5.76 2.47 34.3
Video-ChatGPT 43.5 41.0 40.0 29.0 28.0 5.51 2.65 34.5
VideoChat 32.5 31.5 26.4 24.5 26.0 5.39 2.61 28.2
VideoChat2 54.0 53.5 50.3 30.0 31.0 5.11 3.25 43.8
Video-LLaVA 58.0 47.5 55.8 29.5 34.5 5.38 2.89 45.0
MiniGPT4-Video 57.0 44.5 47.2 27.5 26.5 4.38 2.98 40.5
LLaMA-VID 38.5 36.5 36.0 22.5 25.5 5.31 3.55 31.8
Movie-LLM 23.5 22.5 25.9 25.0 25.0 5.13 3.23 34.4
MA-LMM 48.0 43.0 41.6 27.5 23.5 4.55 2.25 36.7
MovieChat 32.5 29.5 27.4 21.0 21.0 3.60 2.60 26.3
TimeChat 30.5 30.5 31.0 27.5 24.5 4.29 2.92 28.8
Table 7: Quantitative results on MLVU Time-ladder (180s).“M-Avg” is calculated by the average of five multiple-choice tasks. Bold fonts mean the best results.
  Model Needle QA Ego Rea. Plot QA Action Or. Action Co. Sub-Scene Cap. Video Sum. M-Avg
  Otter-I 25.0 25.0 25.0 11.5 25.0 4.17 2.33 22.4
LLaVA-1.6 41.0 38.5 39.6 22.0 26.0 4.14 2.18 33.4
InternVL-1.5 66.5 48.5 58.4 34.0 24.0 5.23 3.12 46.2
Claude-3-Opus 32.0 44.7 42.6 30.2 15.5 3.75 3.13 33.0
Qwen-VL-Max 50.0 43.4 41.1 24.4 14.7 5.49 3.21 34.7
Otter-V 27.0 30.5 23.9 13.0 26.5 4.34 2.51 24.2
mPLUG-Owl-V 29.5 34.5 22.3 23.5 23.0 5.58 2.42 26.6
Video-LLaMA-2 43.0 36.0 40.1 23.0 28.0 5.62 2.45 34.0
Video-ChatGPT 42.5 35.0 27.4 23.5 32.0 5.22 2.56 32.1
VideoChat 28.5 27.5 19.8 25.5 30.0 5.32 2.38 26.3
VideoChat2 47.5 44.5 39.6 27.0 31.0 4.97 2.83 37.9
Video-LLaVA 55.5 47.5 47.7 22.0 35.0 5.25 2.62 41.5
MiniGPT4-Video 53.0 43.5 43.1 31.0 24.5 4.11 2.82 39.0
LLaMA-VID 30.5 39.0 28.4 19.5 30.5 5.39 3.29 29.6
Movie-LLM 26.0 22.5 32.0 22.5 26.0 5.16 3.14 25.8
MA-LMM 41.5 37.0 41.6 27.0 27.5 4.52 2.19 34.9
MovieChat 28.5 25.5 29.4 25.0 23.5 3.25 2.41 26.4
TimeChat 36.5 30.0 31.0 26.5 26.0 4.37 2.66 30.0
Table 8: Quantitative results on MLVU Time-ladder (360s).
  Model Needle QA Ego Rea. Plot QA Action Or. Action Co. Sub-Scene Cap. M-Avg.
  Otter-I 25.0 25.0 25.0 12.5 25.0 4.02 22.5
LLaVA-1.6 37.5 38.5 43.9 25.0 28.5 4.45 34.6
InternVL-1.5 56.0 46.0 55.6 33.0 25.0 5.17 43.1
Claude-3-Opus 27.0 38.5 39.8 35.7 15.0 3.88 31.2
Qwen-VL-Max 43.0 44.8 50.0 17.3 17.9 5.45 34.6
Otter-V 27.5 28.5 22.4 14.0 26.0 4.09 23.7
mPLUG-Owl-V 30.5 36.0 21.4 20.0 28.0 5.29 27.2
Video-LLaMA-2 39.0 39.0 29.6 21.5 25.5 5.62 30.9
Video-ChatGPT 41.0 40.5 28.6 24.5 32.5 5.37 33.4
VideoChat 25.5 34.5 26.0 28.0 30.0 4.91 28.8
VideoChat2 43.0 49.0 41.8 24.5 29.5 5.05 37.6
Video-LLaVA 54.0 49.0 52.0 21.0 29.5 5.21 41.1
MiniGPT4-Video 54.0 39.0 41.8 25.0 28.0 4.13 37.6
LLaMA-VID 32.0 36.5 34.2 27.0 27.0 5.39 31.3
Movie-LLM 26.5 32.5 27.0 21.5 24.5 5.25 26.4
MA-LMM 44.5 41.5 35.7 24.0 28.0 4.72 34.7
MovieChat 27.5 32.5 22.4 24.0 27.0 3.48 26.7
TimeChat 28.0 32.0 23.5 26.0 24.5 4.19 26.8
Table 9: Quantitative results on MLVU Time-ladder (600s).
@@ H.2 Comprehensive evaluation results on MLVU Time-ladder
Appendix J Licensing, Hosting and Maintenance Plan
Author Statement.
We bear all responsibilities for the licensing, distribution, and maintenance of our dataset.
License.
MLVU is under CC-BY-NC-SA 4.0 license.
Hosting.
MLVU can be viewed and downloaded on GitHub at https://github.com/JUNJIE99/MLVU or on Huggingface at https://huggingface.co/datasets/MLVU/MVLU. We assure its long-term preservation for future reference and use. The annotations for questions and answers are provided in the JSON file format, while the raw videos are available in the MP4 format.
For the video files, we do not hold any copyright, and all video rights belong to the video authors. To facilitate user usage, under the premise of a user agreement that the data will be used solely for research purposes and not for commercial use, we provide the download methods for these videos. We have taken measures such as reducing resolution, modifying aspect ratio, and editing to minimize the impact on the original work rights. However, there is still a risk that copyright holders may request the removal of some data. If this happens, we will follow the practice of Movienet [13] to sparsely collect these video frames. This will not significantly impact our data usage, as all our annotations are focused on the visual information in the videos, unrelated to the audio. Moreover, most of the models we evaluate are essentially still frame-based video processing and do not involve audio processing. If retaining video frames is also not allowed, we will still preserve the annotation data and provide metadata for the corresponding videos.
Metadata.
Appendix K Datasheet
K.1 Motivation
For what purpose was the dataset created?
Answer: The creation of MLVU is to facilitate the evaluation and development of MLLM’s long video understanding capabilities. Compared to previous video understanding benchmarks, MLVU possesses a substantial extension of video length, diversified video categories, and diversified evaluation tasks. MLVU is the first comprehensive benchmark for long video understanding.
Who created the dataset (e.g., which team, research group) and on behalf of which entity (e.g., company, institution, organization)?
Answer: The MLVU is created by Junjie Zhou (BUPT, BAAI), Yan Shu (BAAI), Bo Zhao (BAAI), Boya Wu (BAAI), Shitao Xiao (BAAI), Xi Yang (BAAI), Yongping Xiong (BUPT), Bo Zhang (ZJU), Tiejun Huang (PKU, BAAI), and Zheng Liu (BAAI)
Who funded the creation of the dataset?
Answer: Beijing Academy of Artificial Intelligence, Beijing University of Posts and Telecommunications, Peking University, and Zhejiang University.
K.2 Composition
What do the instances that comprise the dataset represent? (e.g., documents, photos, people, countries)
Answer: Each instance in our dataset represents a long video ranging from 3 minutes to 2 hours in duration, a question, a standard answer, and for the multiple-choice task, there are also 4 options. Videos are stored in MP4 file format, while the questions, standard answers, and options are all stored in JSON format files.
How many instances are there in total (of each type, if appropriate)?
Answer: In total, we collect 2593 instances, including 2175 multiple-choice questions and 418 free-form generation questions. The number of distinct videos is 1334, as one video may correspond to multiple questions. The specific distribution of question types and video duration can be found in Figure 1 in the main paper.
Does the dataset contain all possible instances or is it a sample (not necessarily random) of instances from a larger set?
Answer: Only a part of the questions was sampled from existing datasets (the ego reasoning task and the anomaly recognition task). The rest of the questions were newly annotated by us (for example, plot question answering, sub-scene captioning, video summarization, topic reasoning), or they were substantially modified based on existing annotations (needle question answering, action count, action order). During the data collection and annotation process, we ensured that the instances within each task were diverse in terms of videos and questions.
Is there a label or target associated with each instance?
Answer: Yes, for the multiple-choice question type, each instance provides the correct option. For the free-form generation task, each instance comes with a standard caption or summary answer.
Is any information missing from individual instances?
Answer: All instances are complete.
Are relationships between individual instances made explicit (e.g., users’ movie ratings, social network links)?
Answer: Some instances may have the same video but different questions and answers. Each instance clearly indicates the corresponding video, and each video has a unique identifier.
Are there recommended data splits (e.g., training, development/validation, testing)?
Answer: No. Our MLVU is specifically designed for evaluation, with its core objective being to assess the capability of MLLMs to understand long-term video.
Are there any errors, sources of noise, or redundancies in the dataset?
Answer: No.
Is the dataset self-contained, or does it link to or otherwise rely on external resources (e.g., websites, tweets, other datasets)?
Answer: All data will be publicly accessible in the dataset repository. Our annotations will be stored in JSON format. As our collected videos include parts of movies, TV series, documentaries, and cartoons, we have taken measures such as reducing resolution, modifying aspect ratio, and editing to minimize the impact on the original work rights. However, there is still a risk that copyright holders may request removal of some data. If this happens, we will follow the practice of Movienet [13] to sparsely collect these video frames. This will not have much impact on our data usage, as all our annotations are focused on the visual information in the videos, unrelated to the audio. If retaining video frames is also not allowed, we will still keep the annotation data and provide metadata for the corresponding videos.
Does the dataset contain data that might be considered confidential?
Answer: No.
Does the dataset contain data that, if viewed directly, might be offensive, insulting, threatening, or might otherwise cause anxiety?
Answer: No.
K.3 Collection Process
The data collection process is described in Section 3.2 of the main paper and Appendix B, D, and E.
K.4 Uses
Has the dataset been used for any tasks already?
Answer: Yes, MLVU has been used to evaluate the long-video understanding capabilities of as many as 20 different MLLMs. For specific details, please refer to Section 4.
What (other) tasks could the dataset be used for?
Answer: MLVU is primarily used to comprehensively evaluate the long-video understanding capabilities of MLLMs. Since MLVU includes multiple different tasks, it can also be used to individually assess the ability of MLLMs or other video-specific models on a particular task, such as video summarization.
Is there a repository that links to any or all papers or systems that use the dataset?
Answer: No.
Is there anything about the composition of the dataset or the way it was collected and preprocessed/cleaned/labeled that might impact future uses?
Answer: As our collected videos include parts of movies, TV series, documentaries, and cartoons, we have taken measures such as reducing resolution, modifying aspect ratio, and editing to minimize the impact on the original work rights. However, there is still a risk that copyright holders may request removal of some data. If this happens, we will follow the practice of Movienet [13] to sparsely collect these video frames. This will not have much impact on our data usage, as all our annotations are focused on the visual information in the videos, unrelated to the audio. If retaining video frames is also not allowed, we will still keep the annotation data and provide metadata for the corresponding videos.
Are there tasks for which the dataset should not be used?
Answer: The MLVU cannot be used to evaluate the image or video generation capabilities of the MLLM.
K.5 Distribution
Will the dataset be distributed to third parties outside of the entity (e.g., company, institution, organization) on behalf of which the dataset was created?
Answer: Yes. The benchmark is publicly available on the Internet.
How will the dataset will be distributed (e.g., tarball on website, API, GitHub)?
Answer: The benchmark is available on GitHub at https://github.com/JUNJIE99/MLVU or on Huggingface at https://huggingface.co/datasets/MLVU/MVLU.
Will the dataset be distributed under a copyright or other intellectual property (IP) license, and/or under applicable terms of use (ToU)?
Answer: CC-BY-NC-SA 4.0.
Have any third parties imposed IP-based or other restrictions on the data associated with the instances?
Answer: No.
Do any export controls or other regulatory restrictions apply to the dataset or to individual instances?
Answer: No.
K.6 Maintenance
Who will be supporting/hosting/maintaining the dataset?
Answer: The authors will be supporting, hosting, and maintaining the dataset.
How can the owner/curator/manager of the dataset be contacted (e.g., email address)?
Answer: Please contact the official email of our project (mlvubenchmark@gmail.com), or contact the one of the authors (zhoujunjie@bupt.edu.cn; shuyan9812@gmail.com)
Is there an erratum?
Answer: No. We will make announcements if there are any.
Will the dataset be updated (e.g., to correct labeling errors, add new instances, delete instances)?
Answer: Yes. We will post new update in https://github.com/JUNJIE99/MLVU and https://huggingface.co/datasets/MLVU/MVLU if there is any.
If the dataset relates to people, are there applicable limits on the retention of the data associated with the instances (e.g., were individuals in question told that their data would be retained for a fixed period of time and then deleted)?
Answer: People may appear in the reference videos. People may contact us to exclude specific data instances if they appear in the reference videos.
Will older versions of the dataset continue to be supported/hosted/maintained?
Answer: Yes. Old versions will also be hosted in https://github.com/JUNJIE99/MLVU and https://huggingface.co/datasets/MLVU/MVLU.
If others want to extend/augment/build on/contribute to the dataset, is there a mechanism for them to do so?
Answer: If others wish to add data, they can apply to do so provided the data is compliant and reasonable. However, making other modifications based on our dataset is currently not allowed.

### v3

@@ 5 Conclusion
Supplementary Material
@@ Appendix A Overview of Appendix
  • B: Evaluation Results on MLVU Dev Set
  • C: Collecting Details of our Universal Long Video Collection (ULVC)
@@
  • E Detailed Division of Dev and Test Sets in MLVU
  • F: Annotation Details of MLVU
  • G: Details of Baselines and the Evaluation Process.
  • H: Explorations of Video Retrieval Augmented Generation
  • I: More Visualized Examples of MLVU
Appendix B Evaluation Results on MLVU Dev Set
The evaluation results of the baselines on the MLVU dev set are detailed in Table 2. Notably, the multiple-choice questions in the MLVU dev set present four options, whereas the MLVU test set offers six, making the latter more challenging and discriminative.
Appendix C Collecting Details of our Universal Long Video Collection (ULVC)
@@
Specifically, our ULVC includes a diverse set of 986 long videos. This collection features 168 movies from the Movie101 [58] and MovieChat [41] datasets, along with 60 documentaries from MovieChat [41]. It also contains 65 game videos from MineDojo [10], 239 surveillance videos from UCF-Crime [43], and 100 ego-centric videos from Ego4D [15]. Additionally, we independently collected 72 cartoons, 92 TV series, 60 tutorial videos, 60 sports videos, and 70 life records.
It’s important to clarify that the quantity of videos in the ULVC does not directly correspond to the number of videos and questions in our MLVU benchmark, which are 1,730 and 3,102 respectively. For example, a two-hour movie from the ULVC might be utilized in its entirety for the Sub-Scene Captioning task, or it could be segmented into several approximately 10-minute clips for the Video Summarization task, or even used as a background video for synthetic video generation. Moreover, a single video could be annotated with multiple questions simultaneously.
@@ Appendix D Details of the MLVU Time-Ladder
Appendix E Detailed Division of Dev and Test Sets in MLVU
Our MLVU comprises a total of 3,102 questions, divided into a dev set with 2,593 questions and a test set with 509 questions. We present the detailed distribution of questions for each task in Table 1.
Task Dev Test Total
Topic Reasoning 264 91 355
Anomaly Recognition 200 39 239
Video Summarization 217 40 257
Needle QA 355 60 415
Ego Reasoning 352 53 405
Plot QA 539 50 589
Sub-Scene Captioning 201 46 247
Action Order 259 70 329
Action Count 206 60 266
Table 1: Detailed Distribution of Questions in the MLVU Dataset Across Dev and Test Sets for Each Task.
Methods Date Input Holistic Single Detail Multi Detail M-Avg G-Avg
TR AR VS NQA ER PQA SSC AO AC
Full mark 100 100 10 100 100 100 10 100 100 100 10
Random 25 25 25 25 25 25 25 25
Image MLLMs
Otter-I [23] 2023-05 16 frm 25.0 25.0 2.18 25.1 25.0 24.9 4.12 13.1 25.2 23.3 3.15
LLaVA-1.6 [29] 2024-01 16 frm 60.6 41.0 2.11 43.1 38.4 41.0 4.35 25.5 25.7 39.3 3.23
Claude-3-Opus [2] 2024-03 16 frm 67.2 43.5 3.11 21.6 40.2 47.8 3.66 18.2 16.7 36.5 3.39
Qwen-VL-Max [4] 2024-01 16 frm 67.4 63.5 2.71 40.3 40.9 43.3 5.21 25.0 14.8 42.2 3.96
Short Video MLLMs
Otter-V [23] 2023-05 16 frm 24.6 26.0 2.38 28.2 27.6 22.3 4.23 15.1 26.7 24.4 3.31
mPLUG-Owl-V [54] 2023-04 16 frm 28.0 25.0 2.36 24.5 31.8 27.3 5.31 21.2 23.3 25.9 3.84
VideoChat [25] 2023-05 16 frm 33.0 32.0 2.31 27.0 32.1 27.6 5.01 24.3 28.6 29.2 3.66
Video-LLaMA-2 [59] 2024-08 16 frm 54.5 41.5 2.34 39.4 33.5 35.4 5.22 18.5 25.7 35.5 3.78
VideoChat2-HD [26] 2024-06 16 frm 74.6 51.5 2.57 42.0 47.4 43.8 5.04 22.8 29.6 44.5 3.81
Video-LLaVA [28] 2023-11 8 frm 71.6 57.0 2.43 53.2 45.2 48.4 5.25 20.1 35.9 47.3 3.84
ShareGPT4Video [7] 2024-05 16 frm 75.8 51.5 2.52 47.6 43.2 48.4 5.02 34.0 23.3 46.4 3.77
VideoLLaMA2 [9] 2024-06 16 frm 74.6 64.5 2.79 49.9 43.8 45.1 5.18 34.0 27.4 48.5 3.99
Long Video MLLMs
MovieChat [41] 2023-07 2048 frm 29.5 25.0 2.33 24.2 24.7 25.8 3.23 28.6 22.8 25.8 2.78
Movie-LLM [42] 2024-03 1 fps 30.0 29.0 2.88 29.6 24.7 24.1 5.00 20.5 24.8 26.1 3.94
TimeChat [39] 2023-12 96 frm 23.1 27.0 2.54 24.5 28.4 25.8 4.29 24.7 32.0 30.9 3.42
LLaMA-VID [27] 2023-11 1 fps 50.8 34.5 3.22 30.1 32.7 32.5 5.22 23.9 27.8 33.2 4.22
MA-LMM [16] 2024-04 1000 frm 51.9 35.5 2.12 43.1 38.9 35.8 4.80 25.1 24.3 36.4 3.46
MiniGPT4-Video [3] 2024-04 90 frm 70.9 52.5 2.64 49.0 48.6 44.5 4.07 23.2 23.0 44.5 3.36
LongVA [60] 2024-06 256 frm 83.3 58.5 3.39 69.3 50.0 67.2 5.26 38.6 27.2 56.3 4.33
Video-CCAM [11] 2024-08 96 frm 84.9 66.0 2.84 73.2 60.5 66.1 5.19 42.1 38.4 63.1 4.01
Video-XL [40] 2024-09 256 frm 80.3 54.5 3.25 73.8 57.4 67.9 5.02 68.3 40.3 64.9 4.14
GPT-4o [37] 2024-05 0.5 fps 87.4 74.5 4.90 64.8 57.1 65.1 6.69 56.7 46.3 64.6 5.80
[…]
Appendix F Annotation Details of MLVU
F.1 Topic Reasoning (TR).
@@ Appendix F Annotation Details of MLVU
F.4 Needle Question-Answering (NQA).
@@
F.6 Plot Question-Answering (PQA).
@@
F.7 Sub-Scene Captioning (SSC).
In the development process of the SSC task, we employed human annotators to generate both prompts and standard caption data. The specific guidelines provided to annotators are illustrated in Figure 6. Initially, the annotators identified a specific, easily referable sub-scene within a lengthy movie. Subsequently, they crafted a prompt replete with adequate clues to reference this scene, ensuring the uniqueness of these clues throughout the entire film. To prevent any leakage of information, the prompt was designed to exclude any character-specific names or objective hints, instead incorporating rich descriptive details to allude to the plot. Following this, the annotators produced a detailed caption for this sub-scene, and deconstructed the caption into multiple, non-redundant "scoring points" to facilitate quantitative assessment (the details of the evaluation metric can be found in Section G.3). More visualized examples of PQA task can be found in Figure 12.
@@
Appendix G Details of Baselines and the Evaluation Process
G.1 Baselines
In this section, we outline the primary baselines evaluated on our MLVU. For image-based MLLMs, most models lack multi-image inference capabilities. Therefore, we select Otter-I, LLaVA-1.6, and InternVL, which have official multi-image implementations. Additionally, we include two proprietary models—Claude-3-Opus and Qwen-VL-Max—that offer APIs for multi-image inference. For the available models, we determine the maximum input frames based on their LLM context length. Claude and Qwen support a maximum of approximately 20 images, so we choose 16 frames to ensure fair comparisons. Regarding video MLLMs, we use default frame sampling strategies. For example, VideoChat2 uniformly samples 16 frames, while LLaMA-Vid samples 1 frame per second. Specifically, GPT-4o can handle up to approximately 500 images at a resolution of 512×512 pixels. Thus, we select a sampling rate of 0.5 fps to accommodate most of our videos.
G.2 Inference Detatils
@@ G.3 Evaluation Metrics
Appendix H Explorations of Video Retrieval Augmented Generation
@@
We present additional visualizations of our MLVU annotation examples in Figures 10, 11, and 12.